OpenStudy (anonymous):

The concentration (in milligrams per liter) of an antibiotic in the blood is given by the geometric series A + Aekt + Ae2kt + + Ae(n − 1)kt where A is the number of milligrams in one dose of the antibiotic, n is the number of doses, t is the time between doses, and k is a constant that depends on how quickly the body metabolizes the antibiotic. Suppose one dose of an antibiotic increases the blood level of the antibiotic by 0.3 milligram per liter. If the antibiotic is given every 4 hours and k = −0.862 , find the concentration, to the nearest hundredth, of the antibiotic just before t

OpenStudy (anonymous):

The fifth dose

OpenStudy (perl):

the question is a bit ambiguous because it says 'just before the fifth dose' $$ \Large {A + Ae^{kt} + Ae^{2kt} + ...+ Ae^{(n-1)kt} \\ \ \\ = \sum_{i=1}^{n }A e^{(i-1)\cdot kt} = \frac{A\cdot ( e^{knt} - 1)}{e^{kt} - 1 } } $$